# Convergent Primal-Dual Plug-and-Play Image Restoration: A General Algorithm and Applications

### v2

@@ IV-B2 Restoration Performance
To examine how firmly nonexpansiveness of the denoiser affects stability, we compare PnP-PDS (Unstable) with the proposed method. Fig. 4 shows the evolution of PSNR and the update rate $c_{n}$ at each iteration for the deblurring task with $\sigma=0.01$. For the proposed method, the value of $c_{n}$ decreases to approximately $10^{-5}$ and shows a convergent evolution. In contrast, PnP-PDS (Unstable), which uses a denoiser that is not necessarily firmly nonexpansive, shows inconsistent behavior.
@@
In Fig. 5, we provide the visual results for deblurring at $\sigma=0.01$, accompanied by PSNR (left) and SSIM (right). The images obtained by PnP-FBS and TV are overly-smoothed, resulting in a loss of fine detail. For the PnP methods without theoretical convergence guarantees, namely PnP-PDS (Unstable) and DRUNet, strong artifacts were generated because of the unstable behavior of the algorithm (see for example the area above the glove and the texture on the ground in the image, respectively). While RED achieves high PSNR and SSIM, there are some areas where the brightness appears unnaturally altered compared to the original image. The proposed method achieves the highest PSNR and SSIM for this image, presenting the most natural appearance.
Remark 3 (Comparison with Clamping of Denoiser Output).
An alternative approach to enforcing the box constraint is to clamp the denoiser output to [0, 1]01[0,\,1][ 0 , 1 ] after each iteration. This simple post-processing step can help stabilize the algorithm, and we evaluated its effect in the context of PnP-PDS without an explicit box constraint. Empirically, the performance was nearly identical to that of the proposed method across all settings in this section. However, the proposed formulation explicitly incorporates the box constraint into the monotone inclusion formulation. This allows for a clearer characterization of the optimality condition and improves the interpretability of the solution.
@@ IV-C Deblurring/Inpainting under Poisson Noise
Scaling coefficient $\eta$ $50$ $100$ $200$
Deblurring $0.00125$ $0.00125$ $0.00100$
Inpainting $0.00075$ $0.00050$ $0.00050$
TABLE V: The Values of $\lambda$ Used in Each Setting for the Proposed Method.
@@
Scaling coefficient $\eta$ $50$ $100$ $200$ $50$ $100$ $200$
PnP-ADMM [24] $25.16$ $26.28$ $26.55$ $11.45$ $27.87$ $28.10$
PnP-PDS (Unstable) [25] $23.98$ $25.18$ $26.51$ $\mathbf{29.75}$ $\mathbf{31.90}$ $\mathbf{33.43}$
PnP-PDS (w/o a box const.) [26] $\mathbf{25.17}$ $\underline{26.42}$ $\mathbf{27.77}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
BPG [50] $21.76$ $26.36$ $27.45$ $11.84$ $12.06$ $12.18$
RED [44] $25.13$ $26.39$ $27.72$ $29.71$ $31.19$ $32.69$
PnP-PDS (Proposed) $\mathbf{25.17}$ $\mathbf{26.44}$ $\mathbf{27.77}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
TABLE VI: The Average PSNR [dB] Values for Deblurring and Inpainting on the Three Images from Set3[17]. The Best Value is Highlighted in Bold, and the Second Best Value is Underlined.
@@
Scaling coefficient $\eta$ $50$ $100$ $200$ $50$ $100$ $200$
PnP-ADMM [24] $1.692$ $2.224$ $2.898$ $1.203$ $0.6317$ $0.9953$
PnP-PDS (Proposed) $0.6523$ $0.6465$ $0.6529$ $0.4005$ $0.4030$ $0.4039$
TABLE VII: The Average CPU Time [s] per Iteration for Deblurring and Inpainting on the Three Images from Set3[17].
@@
(b) (19.26, 0.2653)19.260.2653(19.26,\,0.2653)( 19.26 , 0.2653 )
@@
Figure 7: Visual results of image restoration on Set3 for deblurring at $\eta=100$. The best results are indicated in bold, and the second-best results are underlined.
@@
Let us now present the experiments on deblurring/inpainting under Poisson noise. We considered three different scaling coefficients $\eta$ for the Poisson noise level, i.e. $50$, $100$, and $200$. Smaller values of $\eta$ correspond to stronger noise, while larger values indicate weaker noise. We compare the performance of the proposed method with other state-of-the-art methods applicable to this scenario, specifically PnP-ADMM [24], PnP-PDS (Unstable) [25], PnP-PDS (w/o a box const.) [26], RED [44], and BPG [50]. For methods whose parameter settings are suggested by the authors, we followed their recommendations. For the others, we performed linear searches over the balancing parameters between the data-fidelity and regularization terms. In particular, the best $\lambda$ in (35) values in the proposed method are summarized in Table V.
The average PSNR values over three images are presented in Table VI. The proposed method shows high restoration performance across all settings and outperforms the other state-of-the-art methods with higher PSNR values.
@@
We investigate the computational efficiency of the proposed method, particularly in comparison with PnP-ADMM. Table VII provides the CPU times per iteration for PnP-ADMM and the proposed method. As discussed in Section III-C2, we need to perform inner iterations for PnP-ADMM, leading to relatively high computational costs. The proposed method reduces CPU times by half compared to PnP-ADMM in all settings, since it solves the problem in (35) without inner iterations.
@@
To evaluate the stability, we present PSNR and the update rate $c_{n}$ at each iteration in the proposed method and PnP-PDS (Unstable) in Fig. 6. For PnP-PDS (Unstable), $c_{n}$ exhibits inconsistently depending on the image. In contrast, the proposed method shows a consistent decrease in $c_{n}$ over iterations, along with stable PSNR values, indicating the stable behavior of the proposed method.
@@
Finally, we present the visual results of the deblurring at $\eta=100$ with PSNR (left) and SSIM (right) in Fig 7. PnP-PDS (Unstable) loses the details of the original images, while PnP-ADMM, RED, and PnP-PDS (w/o a box const.) show slight changes in brightness or fine shapes. Although BPG preserves the structure of the image well, it also produced several visually unpleasant artifacts (e.g., above the man’s shoulder and head). In contrast, the proposed method preserves fine details, resulting in an image that appears visually natural.
@@ IV-D Influence of the Prior on the Proposed Method
Finally, we study how the choice of prior influences the performance of the proposed method. As a representative case, we focus on how the noise level $\sigma_{J}$ assumed during the training of the Gaussian denoiser affects reconstruction performance of deblurring/inpainting under Gaussian noise.
[…]
@@ 2. $(ii)$
$)\displaystyle\mathbf{s}_{n+1}=\rho_{n}((\mathrm{Id}+B)^{-1}(\mathbf{s}_{n})$$)+(1-\rho_{n})\mathbf{s}_{n}.$ (47)
@@
$T_{1}(\mathbf{z}_{n}):=\begin{pmatrix}\gamma_{1}^{-1}A_{J}(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n}\\ -\mathbf{L}\mathbf{x}_{n}+\partial h^{*}(\mathbf{y}_{n})\end{pmatrix},T_{2}(\mathbf{z}_{n}):=\begin{pmatrix}\nabla f(\mathbf{x}_{n})\\ 0\end{pmatrix},$ (50)
@@
where $M_{1}:=P^{-1}\circ T_{1}$ and $M_{2}:=P^{-1}\circ T_{2}$. Since the update in (51) has the same form as (46) with $\gamma=1$, it is sufficient to demonstrate that each assumption in Fact 1 is fulfilled. Hereunder, we define $\kappa$ as follows:
@@
  • We see that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone from [54, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [54, Proposition 20.22]. The operator $\partial h^{*}$ is maximally monotone from [54, Theorem 20.48, Proposition 20.22, and Corollary 16.30]. Therefore, the operator $(\mathbf{x},\mathbf{y})\mapsto\gamma_{1}^{-1}A_{J}(\mathbf{x})\times\partial h^{*}(\mathbf{y})$ is maximally monotone in $\mathcal{Z}_{I}$ from [54, Proposition 20.23]. Furthermore, according to [54, Example 20.35], $(\mathbf{x},\mathbf{y})\mapsto\mathbf{L}^{*}\mathbf{y}\times-\mathbf{L}\mathbf{x}$ is maximally monotone and has a full domain. Therefore, $T_{1}$ becomes an maximally monotone operator on $\mathcal{Z}_{I}$ [54, Corollary 25.5 (i)]. Since $P$ is injective, $M_{1}=P^{-1}\circ T_{1}$ is maximally monotone on $\mathcal{Z}_{P}$.
@@
  • We set $\gamma=1$. Since $10<2\kappa<$ 1 from $(i)$ of Proposition 3.1, the assumption $(i)$ of Fact 1 is fulfilled.
  • The assumption $(ii)$ of Fact 1 is fulfilled from (52).
@@
Therefore, by applying Fact 1, we have that $\mathbf{z}_{n}$ weakly converges to $\mathbf{\hat{z}}\in\mathrm{zer}(T_{1}+T_{2})$ on $\mathcal{Z}_{P}$, which means $\{\mathbf{x}_{n},\mathbf{y}_{n}\}_{n\in\mathbb{N}}$ weakly converges to a solution to (20). ∎
@@
The update in (53) is the same form as Fact 2 by setting $M=M_{1}$, so we need to show that the assumptions of Fact 2 are satisfied. Same as the proof of Proposition 3.1, $M_{1}$ becomes a maximally monotone operator, and from the assumption, $\mathrm{zer}(T_{1}+T_{2})=\mathrm{zer}(T_{1})=\mathrm{zer}(M_{1})$ is nonempty. Furthermore, the assumption on $\rho_{n}$ in Proposition 3.1 is the same as in Fact 2. Therefore, by Fact 2, $\mathbf{z_{n}}$ weakly converges to $\mathrm{zer}(T_{1}+T_{2})$, which means $\{\mathbf{x}_{n},\,\mathbf{y}_{n}\}$ weakly converges to the solution to (20). ∎

### v5

@@ IV-B2 Restoration Performance
To study the impact of enforcing firm nonexpansiveness on the denoiser, we compare the stability of the proposed method with that of PnP-PDS (Unstable) and DRUNet, whose denoisers are not guaranteed to be firmly nonexpansive. Fig. 5 shows the evolution of PSNR and the update rate $c_{n}$ for the deblurring task with $\sigma=0.01$. PnP-PDS (Unstable) exhibits unstable behavior throughout the iterations, and DRUNet shows similar behavior, becoming unstable in the later iterations. In contrast, the proposed method converges stably, with $c_{n}$ decreasing to approximately $10^{-5}$.
@@
In Fig. 6, we provide the visual results for deblurring at $\sigma=0.01$, accompanied by PSNR (left) and SSIM (right). The images obtained by PnP-FBS and TV are overly-smoothed, resulting in a loss of fine detail. For the PnP methods without theoretical convergence guarantees, namely PnP-PDS (Unstable) and DRUNet, strong artifacts are generated because of the unstable behavior of the algorithm (see for example the area above the glove and the texture on the ground in the image, respectively). While RED achieves high PSNR and SSIM, there are some areas where the brightness appears unnaturally altered compared to the original image. The proposed method achieves the highest PSNR and SSIM for this image, presenting the most natural appearance.
@@ IV-C Deblurring/Inpainting under Poisson Noise
Scaling coefficient $\eta$ 11 22 $10$ $50$ $100$ $200$
Deblurring $2.00\times 10^{-3}$ $2.00\times 10^{-3}$ $1.50\times 10^{-3}$ $1.25\times 10^{-3}$ $1.25\times 10^{-3}$ $1.00\times 10^{-3}$
Inpainting $1.25\times 10^{-3}$ $1.25\times 10^{-3}$ $1.00\times 10^{-3}$ $7.50\times 10^{-4}$ $5.00\times 10^{-4}$ $5.00\times 10^{-4}$
TABLE VI: The Values of $\lambda$ Used in Each Setting for the Proposed Method.
@@
Scaling coefficient $\eta$ 11 22 $10$ $50$ $100$ $200$ 11 22 $10$ $50$ $100$ $200$
PnP-ADMM [25] $10.00$ $7.58$ $\mathbf{22.95}$ $25.16$ $26.28$ $26.55$ $7.74$ $9.04$ $25.83$ $11.45$ $27.87$ $28.10$
PnP-PDS (Unstable) [26] $7.07$ $13.86$ $21.06$ $23.98$ $25.18$ $26.51$ $5.38$ $8.25$ $20.29$ $\mathbf{29.75}$ $\mathbf{31.90}$ $\mathbf{33.43}$
PnP-PDS (NoBox) [27] $\underline{18.85}$ $20.08$ $22.85$ $\mathbf{25.17}$ $\underline{26.42}$ $\mathbf{27.77}$ $\underline{18.09}$ $20.47$ $\underline{25.95}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
BPG [51] $5.41$ $6.86$ $16.27$ $21.76$ $26.36$ $27.45$ $6.52$ $7.85$ $10.19$ $11.84$ $12.06$ $12.18$
RED [45] $18.64$ $\underline{20.12}$ $\underline{22.91}$ $25.13$ $26.39$ $27.72$ $18.02$ $\underline{20.75}$ $25.88$ $29.71$ $31.19$ $32.69$
PnP-PDS (Proposed) $\mathbf{19.04}$ $\mathbf{20.19}$ $22.87$ $\mathbf{25.17}$ $\mathbf{26.44}$ $\mathbf{27.77}$ $\mathbf{18.59}$ $\mathbf{21.16}$ $\mathbf{25.96}$ $\mathbf{29.75}$ $\underline{31.21}$ $\underline{32.74}$
TABLE VII: The Average PSNR [dB] Values for Deblurring and Inpainting on the Three Images From Set3[18]. The Best Value is Highlighted in Bold, and the Second Best Value is Underlined.
@@
   PnP-ADMM [25]    $1.525$    $0.5558$
   PnP-PDS (Proposed)    $\mathbf{0.5972}$    $\mathbf{0.3153}$
TABLE VIII: The Average CPU Time [s] per Iteration Across Scaling Coefficients on the Three Images From Set3 [18]. The Better Result is Highlighted in Bold.
@@
(b) (3.471, 0.0257)(3.471,\,0.0257)
Observation
(c) (16.64, 0.6095)(16.64,\,0.6095)
PnP-PDS (NoBox)
[…]
Proposed
Figure 7: Visual results of PnP-PDS (NoBox) and the proposed method for inpainting with $\eta=1$. The higher PSNR and SSIM values are highlighted in bold.
(a) (PSNR, SSIM)
Ground truth
(b) (19.26, 0.2653)(19.26,\,0.2653)
@@
Figure 8: Visual results of deblurring with $\eta=100$. The best results are indicated in bold, and the second-best results are underlined.
@@
Let us now present the experiments on deblurring/inpainting under Poisson noise. To vary the noise intensity, we considered six different values for the scaling coefficients $\eta$ between $1$ and $200$. Smaller values of $\eta$ correspond to stronger noise, while larger values indicate weaker noise. We compare the performance of the proposed method with other methods applicable to this scenario, specifically PnP-ADMM [25], PnP-PDS (Unstable) [26], PnP-PDS (NoBox) [27], RED [45], and BPG [51]. For methods whose parameter settings are suggested by the authors, we followed their recommendations. For the others, we performed a linear search over the balancing parameters between the data-fidelity and regularization terms. In particular, we report the best values of $\lambda$ for the proposed method in Table VI.
The average PSNR values over three images are presented in Table VII. The proposed method shows high restoration performance across all settings and outperforms the other methods with higher PSNR values. Especially, we observe that, at high noise levels (i.e., $\eta\in\{1,\,2\}$), the proposed method outperforms PnP-PDS (NoBox). This indicates that the box constraint not only enhances stability as shown in Section IV-B, but also contributes positively to performance under severely degraded conditions. In Fig. 7, we compare the visual results in such cases. It can be observed that PnP-PDS (NoBox) tends to generate unpleasant artifacts when the noise level becomes high, whereas the proposed method effectively suppresses such artifacts.
Remark 3 (Comparison with Clamping of Denoiser Output).
An alternative approach to enforcing the box constraint is to clamp the denoiser output to [0, 1][0,\,1] after each iteration. This simple post-processing step can help stabilize the algorithm, and using it instead of the box constraint achieves almost the same performance as the proposed method in all settings. On the other hand, the proposed formulation explicitly incorporates the box constraint into the monotone inclusion formulation. This allows for a clearer characterization of the optimality condition and improves the interpretability of the solution.
@@
We investigate the computational efficiency of the proposed method, particularly in comparison with PnP-ADMM. Table VIII provides the CPU times per iteration for PnP-ADMM and the proposed method. As discussed in Section III-C2, we need to perform inner iterations for PnP-ADMM, leading to relatively high computational costs. The proposed method reduces CPU times to approximately half of those of PnP-ADMM in all settings, since it solves the problem in (35) without matrix inversions.
@@
To evaluate stability, we present the PSNR and the update rate $c_{n}$ at each iteration in the proposed method and PnP-PDS (Unstable) in Fig. 9. For PnP-PDS (Unstable), $c_{n}$ behaves unstably depending on the image. In contrast, the proposed method shows a consistent decrease in $c_{n}$ over iterations, along with stable PSNR values.
@@
Fig. 8 shows the visual results of the deblurring at $\eta=100$. PnP-PDS (Unstable) loses the details of the original images, while PnP-ADMM and RED show slight changes in brightness or fine shapes. Although BPG preserves the structure of the image well, it also produced several visually unpleasant artifacts (e.g., above the man’s shoulder and head). In contrast, the proposed method recovers fine details, yielding a visually natural image.
@@ IV-D Influence of the Prior on the Proposed Method
We study how the choice of prior influences the performance of the proposed method. As a representative case, we focus on how the noise level $\sigma_{J}$ assumed during the training of the Gaussian denoiser affects reconstruction performance of deblurring/inpainting under Gaussian noise.
We trained denoisers with four fixed noise levels, specifically $\sigma_{J}\in\{0.0075,\,0.01,\,0.05,\,0.1\}$ where $\sigma_{J}$ denotes the standard deviation of the Gaussian noise added to inputs during denoiser training (i.e., in (17), $\mathbf{x}_{\ell}$ is corrupted with Gaussian noise of standard deviation $\sigma_{J}$). In addition, we trained a mixed-noise-level model, where $\sigma_{J}$ was randomly sampled from the range [0, 0.1][0,\,0.1] for each image. In all cases, the values of $\tau$ and $\xi$ in (17) were manually tuned. For more details, we follow the training procedure described in [18].
@@ V Conclusion
$\mathbf{s}_{n+1}=\rho_{n}((\mathrm{Id}+B)^{-1}(\mathbf{s}_{n})$ $)+(1-\rho_{n})\mathbf{s}_{n}.$ (47)
@@
[…]
@@
where $M_{1}:=P^{-1}\circ T_{1}$ and $M_{2}:=P^{-1}\circ T_{2}$. Since the update in (51) has the same form as (46) with $\gamma=1$, it is sufficient to demonstrate that each assumption in Fact 1 is satisfied. Hereunder, we define $\kappa$ as follows:
@@
  • We see that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone from [55, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [55, Proposition 20.22]. The operator $\partial h^{*}$ is maximally monotone from [55, Theorem 20.48, Proposition 20.22, and Corollary 16.30]. Therefore, the operator $(\mathbf{x},\mathbf{y})\mapsto\gamma_{1}^{-1}A_{J}(\mathbf{x})\times\partial h^{*}(\mathbf{y})$ is maximally monotone in $\mathcal{Z}_{I}$ from [55, Proposition 20.23]. Furthermore, according to [55, Example 20.35], $(\mathbf{x},\mathbf{y})\mapsto\mathbf{L}^{*}\mathbf{y}\times-\mathbf{L}\mathbf{x}$ is maximally monotone and has a full domain. Therefore, $T_{1}$ becomes a maximally monotone operator on $\mathcal{Z}_{I}$ [55, Corollary 25.5 (i)]. Since $P$ is injective, $M_{1}=P^{-1}\circ T_{1}$ is maximally monotone on $\mathcal{Z}_{P}$.
@@
  • We set $\gamma=1$. Since $0<2\kappa<1$ from (i) of Proposition 3.1, the assumption (i) of Fact 1 is satisfied.
  • The assumption (ii) of Fact 1 is satisfied from (52).
@@
Therefore, by applying Fact 1, we have that $\mathbf{z}_{n}$ weakly converges to $\mathbf{\hat{z}}\in\mathrm{zer}(T_{1}+T_{2})$ on $\mathcal{Z}_{P}$, which implies $\{\mathbf{x}_{n},\mathbf{y}_{n}\}_{n\in\mathbb{N}}$ weakly converges to a solution to (20). ∎
@@
The update in (53) is the same form as Fact 2 by setting $B=M_{1}$, so we need to show that the assumptions of Fact 2 are satisfied. Same as the proof of Proposition 3.1, $M_{1}$ becomes a maximally monotone operator, and from the assumption, $\mathrm{zer}(T_{1}+T_{2})=\mathrm{zer}(T_{1})=\mathrm{zer}(M_{1})$ is nonempty. Furthermore, the assumption on $\rho_{n}$ in Proposition 3.1 is the same as in Fact 2. Therefore, by Fact 2, $\mathbf{z_{n}}$ weakly converges to $\mathrm{zer}(T_{1}+T_{2})$, which means $\{\mathbf{x}_{n},\,\mathbf{y}_{n}\}$ weakly converges to the solution to (20). ∎